OpenStudy (anonymous):

Find the midpoint of the line segment whose endpoints are (3, 10) and (7, 8).

OpenStudy (amistre64):

add the points together and divide by 2

OpenStudy (amistre64):

im confident you can do it ;)

OpenStudy (tad1):

use the midpoint formula (x1 +x2)/2 = xmidpoint (Y1 + y2/2) = y midpoint. 3 +7)/2 = 5 ( 10 + 8)/2 = 9 so midpoint isat (5,9)

OpenStudy (anonymous):

Midpoint Formula is the following: [ x1 + x2 y1 + y2 ] [--------- , ----------] [ 2 2 ]

OpenStudy (anonymous):

use midpoint theorem ie x=(x1 +x2)/2 y=(y1+y2)/2

OpenStudy (amistre64):

lol .... yeah, add the points and divide by 2 :) OR first look up a theorum that tells you to add the points and divide by 2 .